OpenStudy (anonymous):

How do I write an inequality in slope intercept form if the line is solid, crosses the x -axis at (0,-2.5) and the y-axis at (5,0), and is shaded up.

OpenStudy (confluxepic):

Welcome to Open Study.

OpenStudy (anonymous):

Thanks

OpenStudy (confluxepic):

If the line is solid then it will be greater than or equal to or less than or equal to.

OpenStudy (anonymous):

Alright, what next?

OpenStudy (confluxepic):

The area where it crosses the x-axis is the x-intercept and y would equal to 0. The area where it crosses the y-axis is the y-intercept and x would equal to 0.

OpenStudy (confluxepic):

First you should find the slope of the line using the two points.

OpenStudy (anonymous):

Using y=mx+b?

OpenStudy (confluxepic):

Then you put it in slope-intercept form.

OpenStudy (anonymous):

How do you find the slope?

OpenStudy (confluxepic):

Slope=(y2-y1)/(x2-x1)

OpenStudy (confluxepic):

Slope=(-2.5-0)/(0-5)

OpenStudy (confluxepic):

Slope=-2.5/-5

OpenStudy (confluxepic):

Slope=1/2

OpenStudy (confluxepic):

You should try to put it in point-slope form.

OpenStudy (confluxepic):

(y-y1)=m(x-x1)

OpenStudy (anonymous):

It has to be in slope intercept form

OpenStudy (confluxepic):

You will convert it later on.

OpenStudy (confluxepic):

(-2.5-0)=1/2(0-5)

OpenStudy (anonymous):

Alright, thanks btw!!!

OpenStudy (anonymous):

What next?

OpenStudy (confluxepic):

-2.5=1/2(-5)+0

OpenStudy (confluxepic):

Make it an inequality by replacing the equal sign with a greater than or equal to sign or a less than or equal to sign.

OpenStudy (anonymous):

Ok